# footnotetext: *Equal Contributionfootnotetext: Corresponding authors

### v1

@@ Abstract
Effective clinical decision-making depends on iterative, multimodal reasoning across diverse sources of evidence. The recent emergence of multimodal reasoning models has significantly transformed the landscape of solving complex tasks. Although such models have achieved notable success in mathematics and science, their application to medical domains remains underexplored. In this work, we propose MedE2, a two-stage post-training pipeline that elicits and then enhances multimodal reasoning for medical domains. In Stage-I, we fine-tune models using 2,000 text-only data samples containing precisely orchestrated reasoning demonstrations to elicit reasoning behaviors. In Stage-II, we further enhance the model’s reasoning capabilities using 1,500 rigorously curated multimodal medical cases, aligning model reasoning outputs with our proposed multimodal medical reasoning preference. Extensive experiments demonstrate the efficacy and reliability of MedE2 in improving the reasoning performance of medical multimodal models. Notably, models trained with MedE2 consistently outperform baselines across multiple medical multimodal benchmarks. Additional validation on larger models and under inference-time scaling further confirms the robustness and practical utility of our approach.
@@ 1 Introduction
Medicine is a multifaceted endeavor involving daily diagnostic and therapeutic activities performed by healthcare professionals. It requires clinicians to process vast amounts of information, including patient medical histories, physical examination findings, and laboratory test results, to diagnose conditions, formulate prognoses, and determine appropriate treatment plans. In many clinical settings, an iterative reasoning process that evaluates multiple possibilities with progressively accumulated clinical information is considered fundamental to medical practice adler2021next ; singh2022five ; croskerry2022advancing . Recent advancements in multimodal large language models (MLLMs), such as OpenAI-o-series openai2024reasoning and Gemini-2.5-Pro Google2025Gemini , have significantly advanced complex task performance. These models employ scaling inference time and emulate reflective cognitive processes, pushing capabilities to unprecedented levels. In light of these advancements, we explore the question: How can we effectively extend the strategy of multimodal reasoning to medical domains?
@@
Figure 1: (a) General MLLMs consistently show superiority over models specifically designed for medical domains across multiple datasets, especially on complex and diverse tasks. (b) Samples that closely mirror real-world clinical scenarios are used to strengthen multimodal reasoning capabilities instead of samples focused primarily on pattern recognition or basic knowledge recall.
Current multimodal reasoning models demonstrating remarkable proficiency are built upon powerful multimodal foundation models bai2025qwen2 ; chen2024internvl that perform well across various general benchmarks yue2024mmmu ; yue2024mmmupro ; meng2025mm ; he2024olympiadbench ; lu2023mathvista . We begin by assessing the capabilities of different models within the medical domain. Such evaluations represent a natural starting point for probing a model’s medical knowledge and reasoning capacities. The results in Figure 1 indicate that current multimodal models specifically tailored to medical tasks chen2024huatuogpt ; li2023llava typically focus only on particular applications, such as visual question answering and medical report generation, and exhibit performance degradation in other medical tasks. In contrast, general multimodal large language models demonstrate superior results. Therefore, we adopt general MLLMs as a foundation for developing specialized medical reasoning models.
@@
Enhancing a model’s capacity for rigorous reasoning depends far more on the quality and domain-specific relevance of its training corpus than on the sheer volume of samples  deepseekai2025deepseekr1incentivizingreasoningcapability . General models predominantly use training samples from mathematical and scientific domains. In contrast, while medical practice encompasses numerous scenarios requiring reasoning, the available data is limited. Unlike previous studies that rely on automatically generated simple question-answer pairs, we implement a rigorous data construction process involving meticulous data collection, cleaning, and expert review. Specifically, we exclude samples that are solvable solely through inherent knowledge or pattern recognition, as shown in Figure 1. Consequently, our carefully curated dataset comprises 2K textual questions sourced from Medical Licensing Examinations jin2021disease and 1.5K multimodal medical questions integrating clinical histories, physical examinations, investigations, and procedures across 12 imaging modalities, including radiology, pathology and optical coherence tomography.
@@
With the dataset in place, we proceed to develop the models’ reasoning capabilities specifically for clinical scenarios. Previous work meng2025mm ; deepseekai2025deepseekr1incentivizingreasoningcapability ; su2025gmai ; li2025limr ; ye2025limo usually employs rule-based reinforcement learning shao2024deepseekmath to advance models’ reasoning capabilities. However, challenges such as normalization of specialized medical terms (e.g., anatomical terminology) and hierarchical relationships (e.g., organ-system classifications) in medical domains make it difficult to verify responses solely through predefined rules. Moreover, methods that utilize neural reward models or reinforcement learning via multiple-choice questions are susceptible to reward hacking denison2024sycophancy . Building upon these, we introduce a novel training pipeline, named MedE2. Our pipeline involves two stages to progressively elicit and enhance the clinical cognitive reasoning capability of the models. In the first stage, we conduct supervised fine-tuning on text-only data, where each sample includes carefully designed reasoning demonstrations to show the behavior that how to utilize existing knowledge to solve complex tasks. Subsequently, we enhance the quality of the generated reasoning process to align with desired patterns. During this stage, we formulate Multimodal Medical Reasoning Preference (MMRP) and utilize Gemini-2.5-Pro to perform rejection sampling on our curated dataset, while employing Direct Preference Optimization (DPO) ziegler2019fine ; ouyang2022training to calibrate and refine the model’s capabilities.
@@
Extensive experiments are conducted to validate the effectiveness of our proposed pipeline, MedE2. Despite being trained only on text-based reasoning data during Stage-I, the models demonstrate significant performance improvements across multiple medical benchmarks, achieving gains of at least 4.45% on MedXpertQA-MM zuo2025medxpertqa and 6.67% on MMMU-Pro-Health yue2024mmmupro . These results highlight an innovative method for eliciting reasoning behaviors in multimodal medical tasks. Further enhancement in Stage-II enables models to exhibit superior reasoning capabilities, reaching performance competitive with several larger-scale models. We also show that improvements from MedE2 generalize well to models with larger parameters, underscoring the scalability of our approach. Moreover, the consistent improvements observed under inference-time scaling further demonstrate the robustness of our post-training recipe.
In summary, our main contributions are as follows:
@@
  • We present MedE2, a two-stage post-training pipeline designed to enhance multimodal reasoning in medical scenarios. The pipeline comprises two critical stages: (1) eliciting reasoning behavior by leveraging strategically developed text-only datasets, and (2) refining the quality of reasoning by incorporating meticulously curated multimodal data.
@@
  • We construct a high-quality dataset of approximately 3.5K samples, comprising both text-based questions with reasoning chains and multimodal questions involving clinical information, thereby establishing a reliable corpus for bootstrapping medical reasoning.
@@
  • We perform comprehensive experiments across diverse benchmarks to validate our post-training pipeline. Experimental results consistently demonstrate that MedE2 significantly enhances model performance and achieves strong generalization across models of varying sizes, while also being robust to inference-time scaling.
@@ 2.1 Medical Multimodal Large Language Models
The development of multimodal large language models has significantly advanced the field of medicine. Current medical multimodal large language models are primarily based on general multimodal models and are further trained on specialized medical datasets. This approach led to the emergence of numerous medical multimodal large models. For example, Med-Flamingo moor2023med builds upon OpenFlamingo-9B awadalla2023openflamingo and continues pretraining with medical image-text data, while Med-PaLM tu2024towards constructs the MultiMedBench dataset and fine-tunes based on PaLM-E driess2023palm . Recent studies continuously use similar strategies and refined training methods, resulting in significant progress, such as LLaVA-Med li2023llava , BioMedGPT zhang2024generalist , MedTrinity-25M xie2024medtrinity , Qilin-Med-VL liu2023qilin , and Med-Gemini saab2024capabilities . These strategies have demonstrated good application results in various medical scenarios, including medical dialogue singhal2025toward , clinical decision support schubert2023evaluating , electronic health record analysis luo2022biogpt , and image report generation li2018hybrid . Given the successful application of reasoning capabilities in scientific tasks, their potential in the medical field is vast; however, effective training strategies to enhance the reasoning abilities of medical models are still lacking. Therefore, our study aims to explore how to enhance the reasoning abilities of multimodal large models in medical tasks.
@@ 2.2 Reasoning Models
Enhancing the reasoning abilities of models remains a key challenge. Early efforts evolved from few-shot prompting to structured paradigms like Chain-of-Thought wei2022chain and ReAct yao2023react , aiming to emulate human-like reasoning. Subsequent work recognized reasoning as an iterative trial, error, and refinement process. This led to the introduction of heuristic search and the process-reward model. Heuristic search, inspired by algorithms such as Monte Carlo Tree Search silver2017mastering , extended single reasoning chains into tree-structured thoughts yao2023tree ; shinn2023reflexion . In parallel, process-reward models framed reasoning as a Markov Decision Process, assigning stepwise rewards to guide the generation of more accurate reasoning chains lightman2023let ; jiao2024learning ; wang2023math . The release of OpenAI’s o1 openai2024reasoning catalyzed further progress. Journey Learning qin2024o1 explored strategies for o1-style slow thinking; o1-coder zhang2024o1 incorporated tree search in code tasks; and STILL-2 min2024imitate distilled long-form reasoning data to expand and refine solution paths. These advancements, culminating in DeepSeek’s results deepseekai2025deepseekr1incentivizingreasoningcapability , highlight the critical role of reinforcement learning in improving reasoning. Diverse reinforcement learning approaches to further bolster reasoning are now being actively explored and investigated wang2024enhancing ; wei2025skywork ; su2025gmai .
Figure 2: Overview of the two-stage post-training recipe MedE2. In Stage-I, text-only data containing reasoning demonstrations is employed to elicit initial reasoning behavior in both textual and multimodal scenarios. In Stage-II, Direct Preference Optimization is applied to multimodal data to further enhance reasoning quality, aligning the model’s outputs more closely with human preferences.
@@ 3 Pipeline
This section presents the core pipeline for adapting multimodal reasoning strategies to the medical domain, as illustrated in Figure 2. To enhance the model’s reasoning capabilities, we first curate a high-quality dataset comprising challenging text and multimodal question-answer pairs (Section 3.1). Based on our pilot studies, we select general MLLMs (e.g., QwenVL2.5 bai2025qwen2 and InternVL3.0 chen2024internvl ) as the base models due to their balanced performance across various medical tasks. Building upon these base models, we propose a two-stage training paradigm that progressively enhances the model’s ability to reason in complex clinical scenarios (Section 3.2 and Section 3.3).
@@ 3.1 Data Collection
We begin by constructing a large-scale question bank sourced from authoritative examinations, top-tier journals, and publicly available datasets. In total, we collect 54K samples, consisting of 37K textual-only and 17K multimodal questions. To ensure high quality and prevent data leakage, a rigorous filtering process was implemented in collaboration with AI experts. Specifically, we employ two models—GPT-4o as the baseline model and Gemini-2.5-Pro as the advanced model—for sample screening. Given that our goal is to explore the potential of multimodal reasoning strategies in solving complex medical tasks, we prioritize filtering out samples lacking sufficient reasoning complexity.
For textual questions, we initially remove samples that GPT-4o could solve directly, indicating they required no meaningful reasoning. The remaining questions are further evaluated by Gemini-2.5-Pro, which attempts each question 4 times. We retain only those samples that the model answers correctly at least once. For multimodal questions, we first assess whether the problem can be solved through mere pattern recognition by prompting GPT-4o to determine the question type. Subsequently, two AI experts conduct independent sampling attempts on the remaining questions. We preserved only those questions correctly answered by at least one expert across all attempts. Ultimately, this process resulted in a dataset consisting of 3.5K samples, including 2K textual and 1.5K multimodal questions.
@@ 3.2 Stage-I: Eliciting Reasoning Ability
To enable models to effectively reason in clinical problem-solving scenarios, it is essential to instruct them on utilizing their existing knowledge base to address complex reasoning tasks. Previous studies ye2025limo ; li2025limr have demonstrated that while supervised fine-tuning with reasoning-specific data can enhance reasoning task performance, reasoning capabilities can be elicited effectively with only a few illustrative examples. Indeed, since much of the relevant knowledge has already been encoded during pre-training, the construction of precisely orchestrated demonstrations of reasoning processes is more critical than merely increasing the volume of training data.
In this stage, we utilize textual-only reasoning demonstrations to elicit sophisticated reasoning capabilities, rather than relying on multimodal reasoning data. This choice is motivated by empirical findings: (1) supervised fine-tuning on multimodal datasets tends to encourage memorization rather than generalization, and (2) entangled multimodal reasoning data can impair the model’s original language reasoning abilities. In contrast, training exclusively on textual data not only enhances reasoning skills but also maintains general visual understanding, albeit with a slight trade-off.
Building upon prior efforts, we adopt a distillation-based method to generate high-quality reasoning demonstrations using our curated dataset. Specifically, we leverage state-of-the-art reasoning models, including o1 openai2024reasoning , DeepSeek R1 deepseekai2025deepseekr1incentivizingreasoningcapability , and QwQ-32B qwq32b , to produce diverse solutions. For open-source models, we directly use their generated reasoning processes. For the proprietary model, such as o1, whose intermediate reasoning steps are not readily accessible, we further prompt the model to expand its summarized outputs into complete reasoning processes. We conduct rigorous evaluations that combine rule-based filtering with human-assisted validation to ensure the quality of generated solutions. We design three criteria: (i) the correctness of the final answer, (ii) logical structure and organization, and (iii) the plausibility and coherence of the reasoning process. In addition, we select a subset of multimodal instances and prompt o1 to elaborate the relevant explanations into detailed reasoning processes, which are filtered using the same criteria and used for comparison.

### v3

@@ Abstract
Effective clinical decision-making depends on iterative, multimodal reasoning across diverse sources of evidence. The recent emergence of multimodal reasoning models has significantly transformed the landscape of solving complex tasks. Although such models have achieved notable success in mathematics and science, their application to medical domains remains underexplored. In this work, we propose MedE2, a two-stage post-training pipeline that elicits and then enhances multimodal reasoning for medical domains. In Stage-I, we fine-tune models using a limited number of text-only data samples containing precisely orchestrated reasoning demonstrations to elicit reasoning behaviors. In Stage-II, we further enhance the model’s reasoning quality using rigorously curated multimodal medical cases, aligning model reasoning outputs with our proposed multimodal medical reasoning preference. Extensive experiments demonstrate the efficacy and reliability of MedE2 in improving the reasoning performance of medical multimodal models. Notably, models trained with MedE2 consistently outperform baselines across multiple medical multimodal benchmarks. Additional validation on larger models and under inference-time scaling further confirms the robustness and practical utility of our approach.
@@ 1 Introduction
Medicine is a multifaceted endeavor. It requires clinicians to process vast amounts of information, including patient medical histories, physical examination findings, and laboratory test results, to diagnose conditions, formulate prognoses, and determine appropriate treatment plans. In many clinical settings, an iterative reasoning process that evaluates multiple possibilities with progressively accumulated clinical information is considered fundamental to medical practice (Adler-Milstein et al., 2021; Singh et al., 2022; Croskerry & Clancy, 2022). Recent advancements in multimodal large language models (MLLMs), such as OpenAI-o-series (OpenAI, 2024) and Gemini-2.5-Pro (Google, 2025), have significantly advanced complex task performance. These models employ scaling inference time and emulate reflective cognitive processes, pushing capabilities to unprecedented levels. In light of these advancements, we explore the question: How can we effectively extend the strategy of multimodal reasoning to medical domains?
@@
Figure 1: (a) Current models’ performance on diverse tasks. (b) Samples that closely mirror real-world clinical scenarios are used to strengthen multimodal reasoning capabilities instead of samples focused primarily on pattern recognition or basic knowledge recall.
We begin by evaluating the capabilities of different models within the medical domain, as such assessments provide a natural starting point for probing a model’s medical knowledge and reasoning abilities. The results shown in the left panel of Figure 1 indicate that current multimodal models (i.e., general-purpose (Bai et al., 2025b; Chen et al., 2024b) and those specifically designed for medical tasks (Chen et al., 2024a; Li et al., 2023) ) demonstrate strong performance on relatively simple visual question answering tasks (Liu et al., 2021; Lau et al., 2018). However, performance declines markedly on more complex tasks (Yue et al., 2024a; b), which require deeper comprehension and advanced reasoning. Since effective reasoning models should be grounded in robust foundation models (Ye et al., 2025), we adopt general-purpose MLLMs as the basis for developing specialized medical reasoning models.
@@
Previous research (DeepSeek-AI, 2025; Ye et al., 2025; Li et al., 2025) primarily relied on carefully curated, challenging datasets from mathematical and scientific domains to incentivize the model’s reasoning capacity. In contrast, although medical practice encompasses numerous scenarios requiring reasoning, the available data remain limited, especially those integrating clinical information. To alleviate this, we adopt a rigorous data construction process involving meticulous data collection, cleaning, and expert review. Unlike earlier studies (Pan et al., 2025; Lai et al., 2025) that relied primarily on simple question–answer pairs solvable through memorized knowledge or pattern recognition (Figure 1), our curated dataset comprises 3K textual questions and 2K multimodal questions spanning 12 imaging modalities, including radiology, pathology, and optical coherence tomography. These samples are sourced from authoritative examinations (Jin et al., 2021) and publicly available case reports published in leading medical journals (PubMed, 2003). Each multimodal case integrates clinical histories, physical examinations, diagnostic investigations, and procedures.
@@
With the dataset in place, we proceed to develop the models’ reasoning capabilities specifically for clinical scenarios. Most of current work (Meng et al., 2025; DeepSeek-AI, 2025; Su et al., 2025; Li et al., 2025) usually employs rule-based reinforcement learning (Shao et al., 2024) to advance models’ reasoning capabilities. However, challenges such as normalization of specialized medical terms (e.g., anatomical terminology) and hierarchical relationships (e.g., organ-system classifications) in medical domains make it difficult to verify responses solely through predefined rules. Moreover, the prevalent strategy that relies on outcome-based rewards is vulnerable to hallucinations during the reasoning process. This poses significant risks in clinical applications, where diagnostic and therapeutic decision-making demands exceptional logical rigor and strong evidence-based justification. Building upon these, we introduce a novel training pipeline, named MedE2. Our pipeline involves two stages to progressively elicit and enhance the clinical cognitive reasoning capability of the models. Instead of cold-start on multimodal data (DeepSeek-AI, 2025), in the first stage, we conduct supervised fine-tuning on text-only data, where each sample includes carefully designed reasoning demonstrations to show how to utilize existing knowledge to solve complex tasks. Subsequently, we enhance the quality of the generated reasoning process to align with desired patterns. During this stage, we formulate Multimodal Medical Reasoning Preference (MMRP) and utilize Gemini-2.5-Pro to perform rejection sampling on our curated dataset, while employing Direct Preference Optimization (DPO) (Ziegler et al., 2019; Ouyang et al., 2022) to calibrate and refine the model’s capabilities.
@@
Extensive experiments are conducted to validate the effectiveness of our proposed pipeline. Despite being trained only on text-based reasoning data during Stage-I, the models demonstrate significant performance improvements across multiple medical benchmarks, achieving gains of at least 4.45% on MedXpertQA-MM (Zuo et al., 2025) and 6.67% on MMMU-Pro-Health (Yue et al., 2024b). These results highlight an innovative method for eliciting reasoning behaviors in multimodal medical tasks. Further enhancement in Stage-II enables models to exhibit superior reasoning capabilities, reaching performance competitive with several larger-scale models. We also show that improvements from MedE2 generalize well to models with larger parameters. Moreover, the consistent improvements observed under inference-time scaling further demonstrate the robustness of our post-training recipe. In summary, our main contributions are as follows:
@@
  • We present MedE2, a two-stage post-training pipeline to enhance multimodal reasoning in medical scenarios: (1) eliciting reasoning behavior by leveraging strategically developed text-only datasets, and (2) refining the reasoning quality by incorporating meticulously curated multimodal data.
@@
  • We construct a high-quality dataset of approximately 5,000 samples, comprising both text-based questions with reasoning chains and multimodal questions involving clinical information, thereby establishing a reliable corpus for bootstrapping medical reasoning.
@@
  • We perform comprehensive experiments across diverse benchmarks. Experimental results consistently demonstrate that MedE2 significantly enhances model performance and achieves strong generalization across models of varying sizes, while also being robust to inference-time scaling.
@@ 2.1 Medical Multimodal Large Language Models
The development of multimodal large language models has significantly advanced the field of medicine. Current medical multimodal large language models are primarily based on general multimodal models and are further trained on specialized medical datasets. This approach led to the emergence of numerous medical multimodal large models. For example, Med-PaLM (Tu et al., 2024) constructs the MultiMedBench dataset and fine-tunes based on PaLM-E (Driess et al., 2023). Recent studies continuously use similar strategies and refined training methods, resulting in significant progress, such as LLaVA-Med (Li et al., 2023), BioMedGPT (Zhang et al., 2024), MedTrinity-25M (Xie et al., 2024), and Med-Gemini (Saab et al., 2024). These strategies have demonstrated good application results in various medical scenarios, including medical dialogue (Ye et al., 2023), clinical decision support (Schubert et al., 2023), electronic health record analysis (Luo et al., 2022), and image report generation (Li et al., 2018). As an endeavor that involves multi-level analysis of details, the reasoning ability is vital for solving medical tasks. In this study, we aim to explore how to incentivize the reasoning abilities of multimodal large models in medical tasks.
@@ 2.2 Reasoning Models
Enhancing the reasoning abilities of models remains a key challenge. Early efforts evolved from few-shot prompting to structured paradigms like Chain-of-Thought (Wei et al., 2022) and ReAct (Yao et al., 2023), aiming to emulate human-like reasoning. Subsequent work recognized reasoning as an iterative trial, error, and refinement process. The release of OpenAI’s o1 (OpenAI, 2024) catalyzed further progress. Journey Learning (Qin et al., 2024) explored strategies for o1-style slow thinking and STILL-2 (Min et al., 2024) distilled long-form reasoning data to expand and refine solution paths. These advancements, culminating in DeepSeek’s results (DeepSeek-AI, 2025), highlight the critical role of reinforcement learning in improving reasoning. Diverse reinforcement learning (RL) approaches to further bolster reasoning are now being actively explored and investigated (Wang et al., 2024; Wei et al., 2025). However, effective training strategies to enhance the reasoning abilities of medical models are still lacking. Previous research has primarily focused on constructing reasoning processes or adapting reinforcement learning (RL) frameworks (Su et al., 2025; Sun et al., 2025; Pan et al., 2025; Lai et al., 2025). In this paper, we propose a novel two-stage post-training recipe that progressively elevates models’ ability to perform fine-grained reasoning in the medical domain.
Figure 2: Overview of the two-stage post-training recipe MedE2. In Stage-I, text-only data containing reasoning demonstrations is employed to elicit initial reasoning behavior. In Stage-II, Direct Preference Optimization is applied to multimodal data to further enhance reasoning quality.
@@ 3 Pipeline
This section presents the core pipeline of MedE2, as illustrated in Figure 2. To enhance the model’s reasoning capabilities, we first curate a high-quality dataset comprising challenging text and multimodal question-answer pairs (Section 3.1). Based on our pilot studies, we select general MLLMs (e.g., QwenVL-2.5 (Bai et al., 2025b) and InternVL-3.0 (Chen et al., 2024b)) as the base models due to their performance across various medical tasks. Building upon these base models, we propose a two-stage training paradigm that progressively enhances the model’s ability to reason in complex clinical scenarios (Section 3.2 and Section 3.3).
@@ 3.1 Data Collection
We start with constructing a large-scale question bank sourced from authoritative examinations, top-tier journals, and publicly available datasets. In total, we collect 54K samples, consisting of 37K textual-only and 17K multimodal questions. To ensure high quality and prevent data leakage, a rigorous filtering process was implemented in collaboration with AI experts.
Given that our goal is to explore the potential of multimodal reasoning strategies in solving complex medical tasks, we prioritize filtering out relatively simple samples. Specifically, we employ two baseline models (i.e., QwenVL-2.5 and InternVL-3.0) to eliminate samples that could be solved directly. For multimodal questions, we further prompt Gemini-2.5-Pro to identify the question type and exclude those classified as pattern-recognition tasks. To ensure that the remaining questions are solvable, AI experts independently attempted each question, with 4 attempts per model (Gemini-2.5-pro (Google, 2025) with DeepSeek-R1 (DeepSeek-AI, 2025) for textual questions and Gemini-2.5-pro with Intern-S1 (Bai et al., 2025a) for multimodal questions). Only samples for which both experts answered correctly at least once were retained. Ultimately, this process resulted in a dataset consisting of 5K samples, including 3K textual and 2K multimodal questions.
@@ 3.2 Stage-I: Eliciting Reasoning Ability
To enable models to effectively reason in clinical problem-solving scenarios, it is essential to instruct them on utilizing their existing knowledge base to address complex reasoning tasks. Previous studies (Ye et al., 2025; Li et al., 2025) have demonstrated that while supervised fine-tuning with reasoning-specific data can enhance reasoning task performance, reasoning capabilities can be elicited effectively with only a few illustrative examples. Indeed, since much of the relevant knowledge has already been encoded during pre-training, the construction of precisely orchestrated demonstrations of reasoning processes is more critical than merely increasing the volume of training data. In this stage, we utilize textual-only reasoning demonstrations to elicit sophisticated reasoning capabilities, rather than relying on multimodal reasoning data. This choice is motivated by empirical findings (Su et al., 2025) that entangled multimodal reasoning data can impair the model’s original language reasoning abilities. In contrast, training exclusively on textual data not only enhances reasoning skills but also maintains general visual understanding, albeit with a slight trade-off.
Building upon prior efforts (Qin et al., 2024; Huang et al., 2024; 2025), we adopt a distillation-based method to generate high-quality reasoning demonstrations using our curated dataset. Specifically, we leverage state-of-the-art reasoning models, including Gemini-2.5-pro, DeepSeek R1, and Qwen3-235B (Team, 2025), to produce diverse solutions. For open-source models, we directly use their generated reasoning processes. For the proprietary model, such as Gemini-2.5-pro, whose intermediate reasoning steps are not readily accessible, we further prompt gpt-OSS-120B (OpenAI, 2025) to expand its summarized outputs into complete reasoning processes. We conduct rigorous evaluations that combine rule-based filtering with human-assisted validation to ensure the quality of generated solutions. We design three criteria: (i) the correctness of the final answer, (ii) logical structure and organization, and (iii) the plausibility and coherence of the reasoning process. Models trained on our curated textual reasoning data, denoted with the “Stage-I” suffix, demonstrate robust reasoning capabilities, not only in tackling textual domains but also when applied to multimodal scenarios.